OpenStudy (experimentx):

Integrate:- \[ \int_0^{\frac \pi 2 } \sin^{2n} (x) dx \]

OpenStudy (jhannybean):

\[\large \int\limits_0^{\frac \pi 2 } \sin^{2n} (x) dx\]so it's readable ^_^

OpenStudy (experimentx):

thanks @Jhannybean

OpenStudy (jhannybean):

This looks like a reduction formula.... \[\large \int\limits_{a}^{b}\sin^n(x)dx\cdots\]

OpenStudy (anonymous):

yes, it looks like a reduction formula.

OpenStudy (jhannybean):

He states that reduction formula won't work here.

OpenStudy (experimentx):

The indefinite integral is ulgy http://www.wolframalpha.com/input/?i=integrate+sin%5E%282n%29x+ according to mathematica the integral evaluates to \[ \int_0^{\frac \pi 2 } \sin^{2n} (x) dx = \frac{\sqrt \pi \Gamma(n + \frac 1 2)}{\Gamma (n +1)} \]

OpenStudy (experimentx):

Woops!! looks like I got this one http://math.stackexchange.com/questions/341402/int-0-pi-2-sin-x1-sqrt2-dx-and-int-0-pi-2-sin-x-sqrt2

OpenStudy (experimentx):

should have realized \[ \sqrt \pi = \Gamma (1/2)\]

OpenStudy (jhannybean):

Lol really? xD explain!!

OpenStudy (experimentx):